OpenStudy (amonoconnor):

I'm stuck on how to work through the reasoning of this one... "Use part one of the Fundamental Theorem of Calculus to find the derivative of the following function: y = integral: [sqrt(t + (sqrt.t))dt], from 0 to tan(x). Any and all help is greatly appreciated!

OpenStudy (amonoconnor):

Here is the function, formatted better: \[y = \int\limits_{0}^{\tan(x)}\sqrt{(t + \sqrt{t})}*dt\]

OpenStudy (irishboy123):

just process the derivative \(\color{#0cbb34}{\text{Originally Posted by}}\) @amonoconnor Here is the function, formatted better: \[y = \int\limits_{0}^{\tan(x)}\sqrt{(t + \sqrt{t})}*dt\] \(\color{#0cbb34}{\text{End of Quote}}\)

OpenStudy (amonoconnor):

...what do you mean "just process the derivative"? I set u = tan(x), but what's the inner function?

OpenStudy (irishboy123):

hey! your task is to find the derivative of the underlying function. **find the derivative of the following function: y = integral: [sqrt(t + (sqrt.t))dt], from 0 to tan(x).**